DISK-OUTFLOW SYSTEM AROUND A MASSIVE PROTOSTAR

\underline{KEE-TAE KIM}, Department of Astronomy, University of Illinois, 1002 W. Grenn St., Urb ana, IL 61801; E. CHURCHWELL, Department of Astronomy, University of Wisconsin at Madison, 475 North Charter St., Madison, WI 53705.

There are two major competing models for massive star formation: accretion via disks and coalescence of low-mass (proto)stars. The presence of well-defined disks around massive protostars may be the most convincing observational evidence to support the accretion model. We observed a massive protostar in the H^{13}CO^{+} and H^{13}CN (1−0) lines and 3-mm continuum at 4′′ resolution using the BIMA array, and detected a well-elongated dense condensation in the NE-SW direction. Dust continuum emission peaks around its center where a bright (∼4×10^{4} L_{⊙}) IRAS point source is located. The condensation crosses the associated outflow axis at about 50^{∘}. A large (∼15 km s^{−1} pc^{−1}) velocity gradient exists along the major axis, suggesting rotation. The condensation has a size of 50000 AU and a mass of 400 M_{⊙}, which are very similar to those of the circumstellar disk recently detected around NGC 7538S. Therefore, this structure seems to be a circumstellar disk around a massive protostar, although further studies are required to confirm this.